Last year there were 879 challenges made to referee calls in professional tennis singles play. Among those challenges, 231 were upheld. Historically,25% of calls are upheld. Find the probable that among the 879 challenges last year 231 or more were upheld
1
Expert's answer
2019-10-17T09:17:21-0400

Let "X=" the number of upheld challenges: "X\sim B(n,p)."

Given that "p=0.25,n=879." Then


"np=879(0.25)=219.75\geq10""n(1-p)=879(1-0.25)=659.25\geq10"


We can use a Normal Distribution to approximate a Binomial Distribution


"X^*\sim N(\mu=np,\sigma^2=np(1-p))""\mu=np=879(0.25)=219.75""\sigma^2=np(1-p)=879(0.25)(1-0.25)=164.8125"

"X^*\sim N(219.75,164.8125)"

Then


"Z={X^*-\mu \over \sigma}\sim N(0,1)"

Find the probability that among the 879 challenges last year 231 or more were upheld


"P(X^*\geq231)=P\big(Z\geq{231-219.75 \over\sqrt{164.8125}}\big)="

"=1-P\big(Z<{231-219.75 \over\sqrt{164.8125}}\big)\approx1-0.80956906="

"=0.19043094"